### v2

# A Joint Learning Framework for Bridging Defect Prediction and Interpretation

Abstract
Understanding why defect predictors classify software components as defective or clean is essential for software engineers that helps identify the root causes of defects and develop actionable bug-fixing plans. Existing solutions employ various explainable artificial intelligence (XAI) methods to clarify the decision-making processes of defect predictors. However, these post-hoc explanation techniques have two main limitations: 1. the interpretation results do not accurately reflect the model’s decision logic, and 2. they do not contribute to improving the performance of defect predictors. To address these limitations, we treat defect prediction and its corresponding interpretation as two distinct but closely related tasks, proposing a joint learning framework that trains the predictor and its interpreter simultaneously. The novelty of our approach lies in two main aspects: 1. we design a feedback loop that transfers decision logic from the predictor to the interpreter, ensuring a high degree of conciseness for both components. 2. we incorporate interpretation results as a penalty term in the loss function of the joint learning framework. This not only enhances the accuracy of the predictor but also proposes a stronger constraint on the reliability of the interpreter. We validate our method against several existing explainable software defect predictors across multiple datasets. The results demonstrate its effectiveness. The source code of our method is available at: https://github.com/BugPredictor/software-defect-prediction.git
Index Terms:
software defect prediction, interpretability, knowledge distillation, joint learning framework
I Introduction
Software defect prediction (SDP) has attracted significant attention in software engineering for over five decades [1, 2]. The primary goal of this research is to identify defective code early in the software development process, helping practitioners prioritize quality assurance tasks, especially when resources are limited or deadlines are tight [3].
In recent years, inspired by the success of machine learning (ML) and deep learning (DL) in various disciplines, researchers have introduced different ML/DL algorithms into SDP. Despite significant improvements in the performance of SDP models, many practitioners remain hesitant to integrate SDP techniques into practical applications [4]. Such a situation is primarily due to the lack of interpretability of these models. In practical scenarios, practitioners often ask: What is the reason for predicting a code fragment as buggy? How do different metrics influence the prediction? These questions underscore an urgent requirement for explainability in SDP [5, 6, 7].
To address the interpretation issue, researchers have proposed various explainable SDP models that integrate different explainable artificial intelligence (XAI) techniques with SDP. Existing explainable SDP models are generally categorized into two groups [8]: global and local approaches. Global approaches focus on explaining the overall behavior of a defect predictor, offering insights into how the model makes decisions across all instances [9]. In contrast, local approaches provide clear evidence for why a given instance is predicted as buggy or clean. Compared to global approaches, local approaches focus on identifying the root cause of specific defects and provide actionable guidance for practitioners to address key questions, such as why a specific prediction was made and how to mitigate associated risks [10]. As a result, local interpretation techniques are generally considered more favorable in practice [5, 11].
However, the task properties of interpretation and defect prediction are quite different, making explainable SDP a challenging problem. On one hand, SDP relies on high-complexity ML/DL models to capture intricate relationships among software metrics and defect patterns. These models are often considered as black boxes, with decision-making processes that are difficult to interpret. On the other hand, XAI methods favor simpler models that provide a clear, human-understandable logical reasoning path. Bridging these two competing objectives—complexity and interpretability—remains a significant challenge. Most existing solutions adopt a post-hoc explanation strategy, in which an SDP model is first trained and subsequently interpreted by XAI techniques [12, 13]. Although this strategy offers a degree of interpretability, it presents two main challenges: 1. the interpretations often fail to reliably reflect the model’s actual decision logic, and 2. they do not contribute to improving the predictive performance of defect models.
[The interpretation for the SMOTE-based predictor] Refer to caption [The interpretation for the RUS-based predictor] Refer to caption
(a) (a) The interpretations by different sampling techniques
[The first interpretation] Refer to caption [The second interpretation] Refer to caption
(b) (b) Two interpretations produced by LIME
Figure 1: An example of unreliable interpretations.
To better illustrate the challenges associated with local approaches, we examine the interpretation of ”AbstractCommand.java” from the Apache ActiveMQ project as an example. The explanations generated by LIME using different data sampling techniques (i.e., SMOTE and RUS) are shown in Fig. 1 LABEL:sub@fig_1:LIME_RUS, and the explanations generated by executing LIME twice are shown in Fig. 1 LABEL:sub@fig1:LIME_two. The features highlighted in red contribute to the prediction of buggy code, while those highlighted in green contribute to the prediction of clean code. From the Fig. 1 LABEL:sub@fig_1:LIME_RUS, we observe that the explanations vary significantly. Among the ten features selected by SMOTE-based LIME, only two are also selected by RUS-based LIME. Fig. 1 LABEL:sub@fig1:LIME_two further illustrates the issue of poor reliability. In the first round of explanation, seven features are identified as correlated with buggy code, whereas in the second round, this number increases to eight.
To address this issue, we treat defect prediction and the corresponding interpretation as two distinct yet closely related tasks, proposing a joint learning framework in this paper. Specifically, we design multiple feedback loops between the interpreter and predictor to simultaneously achieve reliable interpretation and high prediction accuracy. The main contributions are as follows:
  • To enhance the reliability of interpretation, we introduce two novel loss functions: output fidelity loss and feature fidelity loss. This strategy establishes a knowledge transfer loop between the interpreter and predictor, ensuring both a high level of conciseness in decision-making logic and feature engineering power between these two components.
  • To improve the accuracy of the predictor, we incorporate the interpretation results as a penalty term in the loss function of the joint learning framework. This strategy enables the predictor to capture more discriminative features guided by the interpretation results while simultaneously imposing stronger constraints on the fidelity of the interpreter to the predictor.
  • Extensive experiments are conducted on twenty datasets with an in-depth analysis. The results demonstrate that the joint learning framework improves both the reliability of interpretation and the accuracy of the predictor.
The remainder of paper is organized as follows: Section II provides an overview of the existing research on SDP and interpretation methods used in SDP. Section III describes our proposed method in detail, while Section IV presents the experimental setup and results. Section V discusses the threats to validity. Finally, Section VI presents our conclusions and outlines directions for future work.
II Related Work
II-A Software defect prediction
Constructing an SDP model involves three steps. The first step is to transform software metrics, such as cohesion, coupling, complexity, into numerical representations [14]. The second step labels code fragments as buggy or clean based on post-release defects. Finally, leverage the labeled code fragments to train an ML/DL-based classifier for identifying future defect-prone code.
Over the past fifty years, numerous ML algorithms, including fuzzy self-organizing maps, K-means, SVM, Random Forest, and logistic regression, have been extensively employed in SDP research [15, 16, 17, 18, 19, 20, 21]. However, software metrics often exhibit correlations and redundancies [22], and traditional ML algorithms struggle to capture the complex dependencies among them [23]. To address this issue, various DL models have been introduced in SDP. One of the earliest explorations, Deeper [24], employed a deep belief network to identify discriminative information in input data. Experimental results showed that Deeper identified 32.22% more defects than many traditional ML-based approaches. Inspired by the Deeper, researchers have introduced a sophisticated DL algorithms into SDP to enhance performance and broaden application scenarios. For example, [25, 26] leaverage convolutional neural networks (CNNs) to build defect prediction models, achieving better performance than other DL-based approaches. Additionally, Qiao et al. employed a DL model to predict the number of defects [27]. Experimental results showed a 14% reduction in mean squared error and an 8% increase in squared correlation coefficient.
To ensure generalizability, we employ DP-CNN, proposed in [26], as the predictor and evaluate the accuracy differences before and after integrating it into the joint learning framework.
II-B Explainable defect prediction
Efforts in explainable SDP can be broadly divided into two main categories [8]: global and local approaches. The global approaches provide a high-level view of how inputs features influence the model’s predictions. Many ML algorithms (e.g., decision trees, logistic regression) and statistical methods (e.g., ANOVA, variable importance) fall into this category[28, 29]. However, global approaches cannot offer detailed interpretations for individual code fragment, and their simple structure often fails to ensure optimal predictive accuracy. As a result, many practical applications have adopted local methods instead.
Local interpretation methods such as LIME [12] and BreakDown [30], primarily focus on interpreting specific code fragments. However, recent studies have reported that local techniques often lack reliability under various conditions [31, 32, 11]. Firstly, interpretations can be inconsistent when different data sampling techniques are used or when the same interpreter is executed multiple times. Local interpretation methods typically employ an interpreter to approximate the predictor’s behavior within a local area of the given instance. The local area is generated by the sampling algorithm. However, the randomness of the sampling process makes it difficult to maintain consistency across local areas—regardless of whether the same or different sampling methods are used [11]. As a result, the same code fragment may yield distinct interpretations[32]. Secondly, interpretations tend to be oversimplified. Local interpretation methods usually employ ML algorithms with simple structure as interpreters. For example, LIME utilizes a linear regression model as the interpreter. The interpretation of SDP is oversimplified as a linear transformation between input metrics and prediction outcomes. Recent study [32] reported that attempting to comprehend an intricate model by employing a simple model might be overly optimistic. Such interpretations often fail to capture the underlying decision-making logic of the predictor.
II-C Knowledge distillation
Knowledge distillation (KD) is a model compression technique that transfers knowledge from a teacher model (e.g., DL models) to a student model (e.g., a shallow neural network) [33, 34]. It is considered a viable solution for enhancing the reliability of interpretation due to the knowledge transfer mechanism between the teacher and student models. For instance, [35] utilized distilled knowledge to identify disease patterns, while [36] distilled dynamic pricing knowledge from a complex black-box DL model. In this paper, we treat the predictor and the interpreter as the teacher and student models, respectively, and design knowledge transfer channels based on the KD principle. The key distinction of our approach lies in the collaborative training mechanism: unlike existing KD methods, which keep the parameters of the teacher model fixed, our method trains both the interpreter and the predictor in a collaborative manner.
III Proposed Method
Fig. 2 provides an overview of our proposed approach, which consists of four modules: metric selection, predictor and interpreter design, joint learning framework, and local interpretation.
Figure 2: The framework of our approach.
III-A Metric selection
We selected sixteen metrics for our method based on a review of existing literature [14, 37, 38]. The details of these metrics are presented in Table I. The selection was guided by two criterias: (1) expressiveness: The metrics needed to be sufficiently expressive to capture subtle variations in code quality, structure, and complexity. (2) effectiveness: Each selected metric has been demonstrated its utility in SDP tasks [17, 24, 39, 40]. According to these criteria, we constructed a dataset that encompasses a broad spectrum of code attributes, ranging from complexity measures to process-related characteristics.
TABLE I: Metrics suite
Feature Description
WMC Weighted methods per class
DIT Depth of Inheritance Tree
NOC Number of Children
CBO Coupling between object classes
RFC Response for a Class
LCOM Lack of cohesion in methods
NPM Number of Public Methods
DAM Data Access Metric
MOA Measure of Aggregation
MFA Measure of Functional Abstraction
CAM Cohesion Among Methods of Class
IC Inheritance Coupling
CBM Coupling Between Methods
AMC Average Method Complexity
LOC Lines of Code
$C$ McCabe’s cyclomatic complexity
III-B Defect prediction and interpretation model
In this section, we provide the detailed information about the predictor and interpreter.
III-B1 Defect predictor
In this paper, we chose a DL-based model, DP-CNN [26] as the predictor. It consists of three main components: a convolution layer, a max pooling layer, and a classification layer. To further improve predictive accuracy, we extend the original model by integrating an attention mechanism. The overall architecture of DP-CNN is illustrated in Fig. 3.
Figure 3: Model Architecture of DP-CNN.
The DP-CNN model predicts defect-proneness through the following process: Selected metrics are first fed into the convolutional layer to capture correlations among features. The output is then passed through the max pooling layer to highlight key features while reducing redundancy. Next, the attention layer focus on critical defect patterns by dynamically assigning weights. Finally, a fully connected network followed by softmax classifier generates the prediction. We choose the CNN-based model as the predictor for three key reasons: (1) Sparse connectivity enables a wider receptive field to capture non-linear correlations. (2) Weight sharing ensures positional invariance in feature extraction [26]. (3) Max pooling enhances robustness by preserving significant features, while the attention layer further refines feature importance.
III-B2 Interpretation model
Following the KD principle, we employ the soft decision tree (SDT) as the interpreter [41]. The SDT is a variant of the fuzzy decision tree, sharing a similar structure with the ordinary decision tree but differing in the definition of nodes. In SDT, nodes are defined as neurons with learnable weights $W$ and biases $b$. The output of each node is defined as $P_{i}(x_{i})=\sigma(x_{i}W_{i}+b_{i})$, which determines the probability of transitioning to the right subtree. The $\sigma(\cdot)$ denotes the activation function and $x_{i}$ is the input of the $i$-th node. The output at the $l$-th leaf node is defined as:
$Q_{k}^{l}=\frac{{\rm exp}(\phi_{k}^{l})}{\sum_{k}{\rm exp}(\phi_{k}^{l})}$ (1)
where $Q_{k}^{l}$ is the probability at the $l$-th leaf of type $k$ defect, and $\phi_{k}^{l}$ is the learned feature at that leaf.
The reasons we use SDT as an interpreter are as follows: Owing to its neural-like node structure, SDT exhibits comparable feature engineering capabilities to CNN, enabling it to simulate the decision-making knowledge of CNN. Furthermore, the complexity of SDT has been reduced, making the decision-making process more comprehensible.
III-C Joint learning framework
The architecture of the joint learning framework is presented in Fig. 4, each feedback loop corresponds to a loss function. Given training data set $D=\{\left(x_{i},y_{i}\right)_{i=1}^{N}\}$, $x_{i}\in R^{n}$, $y_{i}\in\{0,1\}$, the predictor $f$ and interpreter $g$, the loss function for joint learning framework is defined as follows:
Figure 4: The architecture of joint learning framework.
$L=\alpha L_{pred}(f,D)+\beta L_{pred}(g,D)+L_{int}(f,g,D)$ (2)
where the hyperparameters $\alpha$ and $\beta$ denote the relative importances of the subloss to the the overall loss.
The original loss function of KD consists of two parts: the distillation loss $L_{soft}$ (which is equivalent to $L_{int}(f,g,D)$ in Eq. (2)) and the student loss $L_{hard}$ (which is equivalent to $L_{pred}(g,D)$ in Eq. (2)). $L_{soft}$ quantifies the difference between the teacher model and the student model, which serves as a feedback loop for transferring knowledge from the teacher model to the student model. $L_{hard}$ evaluates the difference between the student model’s predictions and the ground truth.
In this paper, we introduce two key modifications to the original KD loss function: (1) added the output fidelity loss $L_{pred}(f,D)$, and (2) revised the $L_{soft}$. $L_{pred}(f,D)$ shares a similar definition with $L_{pred}(g,D)$ indicating the consistency of the predictor’s output to the ground truth.
$L_{pred}(f,D)=\sum_{i=1}^{|D|}\left(f(x_{i})-y_{i}\right)^{2}$ (3)
$L_{pred}(g,D)=\sum_{i=1}^{|D|}\left(g(x_{i})-y_{i}\right)^{2}$ (4)
Introducing output fidelity loss here can enhance both prediction accuracy and interpreter fidelity simultaneously. Within the KD framework, predictor and interpreter are mutually dependent. Minimizing the output fidelity loss will encourage the predictor to achieve heigher accuracy. Moreover, improving the accuracy of the predictor is equivalent to indirectly enforcing the interpreter to provide a more accurate explanation of the predictor’s decision logic. Therefore, incorporating the output fidelity loss can boost the performance of the predictor but also yield a more robust and insightful knowledge based on the KD principle.
The revised the loss function $L_{soft}$ (denoted as $L_{int}(f,g,D)$ in Eq. (2)) is presented in Eq. (5). It involves two parts, $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$. $L_{of}(f,g,D)$ is equivalent to the distillation loss in the original KD framework. The feature fidelity loss $L_{fef}(f,g,D)$ evaluates the similarity between the feature maps generated by predictor and interpreter. Minimizing $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$ can enforce the interpreter to achieve similar feature engineering power of predictor.
$L_{int}(f,g,D)=\lambda L_{of}(f,g,D)+\gamma L_{fef}(f,g,D)$ (5)
where $\lambda$ and $\gamma$ are hyperparameters. $L_{of}(f,g,D)$ shares a similar definition with $L_{soft}(f,g,D)$.
$L_{of}(f,g,T)=-\sum_{i=1}^{|D|}q_{i}^{T}\log(p_{i}^{T})$ (6)
where $q_{i}^{T}=\frac{{\rm exp}(\nicefrac{{z_{i}}}{{T}})}{\sum_{k}^{N}{\rm exp}(\nicefrac{{z_{k}}}{{T}})}$ and $p_{i}^{T}=\frac{{\rm exp}(\nicefrac{{v_{i}}}{{T}})}{\sum_{k}^{N}{\rm exp}(\nicefrac{{v_{k}}}{{T}})}$ are the classification probability of instance $x_{i}$ generated by $g$ and $f$, $z_{i}$ and $v_{i}$ are the logits of $x_{i}$. $T$ is the temperature hyperparameter [41]. Furthermore, the feature fidelity loss $L_{fef}$ is defined as follows:
$L_{fef}=\sum_{i}^{|D|}\left(d(z_{i})-v_{i}\right)^{2}$ (7)
where function $d(\cdot)$ adjusts $z_{i}$ to the same dimension of $v_{i}$.
To further demonstrate the proposed joint learning framework, we use the PC2 dataset as an example. Each data point in PC2 consists of twenty software metrics and one label. According to Eqs. (3, 4), DP-CNN and SDT iteratively optimize their parameters to minimize discrepancies between predictions and actual labels. Crucially, improving prediction accuracy of DP-CNN inherently enforces the SDT to provide a more precise explanation of the predictor’s decision-making process, which indirectly improve interpreter’s reliability. Meanwhile, according to Eq. (6), the features generated by DP-CNN and SDT should be similar, ensuring that both models have comparable decision logic, thereby enhancing the interpretability of SDT. Additionally, Eq. (7) ensures consistency between SDT’s selected features and the input metrics, which strengthens the reliability of the interpreter. As presented in Fig. 5, the decision logic of DP-CNN is interpretered as a set of “IF-THEN” paths in the SDT, which is similar to the decision tree doess [41].
Figure 5: The interpretations of the proposed method on PC1 dataset.
III-D Local interpretation
In this section, we focus on the interpretability for an instance. Given a metric set $A=\{a_{1},a_{2},...a_{d}\}$, the sensitivity of SDT for an instance $x$ on the $a_{i}$ is defined as follows.
$S(a_{i})=\frac{\Delta g_{i}(x)}{\Delta a_{i}}$ (8)
$\Delta g_{i}(x)=\lvert g(x)-g(x+{\rm One}(a_{i})\Delta a_{i})\rvert$ is the changes of the output of interpreter aroused by the perturbation of the metric $a_{i}$. $\Delta a_{i}$ is the standard variance of $a_{i}$. ${\rm One}(a_{i})$ is the onehot encoder of $i$-the metric.
$\Delta a_{i}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(a_{i}-\bar{a})^{2}}$ (9)
$\bar{a}_{i}=\frac{1}{N}\sum_{j}^{N}a_{j}^{i}$ (10)
where $N$ is the number of training data, and $a_{j}^{i}$ is the $i$-th metric of $x_{j}$. According to Eq. (8), we interpret the correlation between the instance and predictive result as the sensitivity to feature perturbation. Given an instance $x_{i}$, if a metric has a greater sensitivity, this means the metric has a stronger influence on the prediction result.
Since the $\Delta a_{i}$ is a constant and the parameters of interpreter are fixed after training, the sensitivity $S$ can provide a stable and consistent interpretation for any instance. The algorithm for performing local interpretation is presented below:
Algorithm 1 Local Interpretation
0: interpreter $g$, instance $x$, metric set $A=\{a_{1},a_{2},\dots,a_{d}\}$, and hyperparameters: $\lambda,\gamma$.
0: Sensitivity index SISI
1:for $a_{i}\in A$ do
2:  $\text{Calculate }\bar{a}_{i}$ Eq. (10)
3:  $\text{Calculate }\Delta a_{i}$ Eq. (9)
4:  $\text{Calculate }\Delta g_{i}(x)$
5:  $S=\cup\ S(a_{i})$ Eq. (8)
6:end for
[…]
8:return SISI
The loop structure (lines 2 to 5 in the algorithm) computes the sensitivity of each metric and generates the sensitivity set $S$. Line 7 rearranges the sensitivity set $S$ in descending order and puts the index of $S$ to $SI$.
IV Experiment
To objectively evaluate the performance of our method, we conducted an experiment to quantify the performance of the proposed method and the baselines on twenty datasets, aiming to answer the following research questions:
RQ1: What improvements does the proposed method offer for the consistency of interpretations?
As mentioned in Section II, many local methods tend to yield inconsistent interpretations, which greatly affects the reliability of explainable SDP. To evaluate the consistency of interpretation, we adopt the Coincidence Degree (CD) in RQ1 [42]. Given an interpreter $g$, the CD of an instance $x_{i}$ is defined as follows :
$CD(x_{i})=\frac{1}{|M|}\left|\bigcap_{j=1}^{|M|}g_{j}(x_{i})\right|$ (11)
where $\bigcap_{j=1}^{|M|}{g_{j}(x_{i})}$ is the intersection of the multiple interpretations for one instance, $|M|$ represents the number of interpretations or the number of data sampling techniques used. $CD-k\%$ indicates the coincidence of the top $k$ metrics. A greater $CD$ indicates greater consistency.
Furthermore, another issue related to the consistency of interpretation is that the predictor and interpreter may generate different predictive results for the same instance. To evaluate the consistency between predictor and interpreter, we introduce Fidelity of Interpretation (FI) [42, 43, 44] here. Given dataset $D$, predictor $f$ and interpreter $g$, the fidelity is defined as:
$FI=\frac{1}{|D|}\sum_{(x_{i},y_{i})\in D}\mathbb{I}\left(f(x_{i}),g(x_{i})\right)$ (12)
$\mathbb{I}(\cdot)$ is an indicator function, which generates 1 when $f(x_{i})=g(x_{i})$, 0 otherwise.
In addition to FI, we also argue that the prediction produced by an interpreter should also align with the ground truth. The alignment indicates how much correct decision knowledge is captured from the predictor. In this paper, we use the Accuracy of Interpreter (AI) to measure this alignment. Given the dataset $D$ and interpreter $g$, we define the accuracy as:
$AI=\frac{1}{|D|}\sum_{(x_{i},y_{i})\in D}\mathbb{I}\left(g(x_{i}),y_{i}\right)$ (13)
where $\left|D\right|$ denotes the cardinality of $D$, $g\left(x_{i}\right)$ is the prediction made by the interpreter for the instance $x_{i}$.
RQ2: What performance improvements does the joint-learning framework achieve?
There are two goals in validating the predictor trained by the joint learning framework (denoted as Joint_CNN): (1) to assess the positive impact of the joint learning framework on predictor’s performance, and (2) to determine whether its predictive performance outperforms other notable predictors. For the first goal, we evaluate model performance using five key indicators: F-measure, AUC, Percent of Perfect Cleans (PPC), Percent of Non-Perfect Cleans (PNPC), and False Omission Rate (FOR). Although F-measure and AUC provide an overall assessment of predictive performance, they can not fully capture the effectiveness in practical SDP scenarios. To address this gap, we introduce PPC, PNPC, and FOR. For the second goal, we compare the F-measure and AUC of Joint_CNN with existing explainable and non-explainable SDP models. PPC, PNPC and FOR are defined as follows:
$PPC=\frac{\left|TN\right|}{\left|D\right|}$ (14)
$PNPC=\frac{\left|TP\right|+\left|FP\right|+\left|FN\right|}{\left|D\right|}$ (15)
$FOR=\frac{\left|FN\right|}{\left|TN\right|+\left|FN\right|}$ (16)
where $\left|TP\right|,\left|FP\right|,\left|FN\right|,\left|FN\right|$ represent the number of true positives, false positives, true negative and false negatives.
The F-measure assesses the balance between precision and recall.
$F\text{-measure}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (17)
The AUC is used to evaluate the performance of binary classification models by measuring the area under its Receiver Operating Characteristic (ROC) curve.
RQ3: what’s the performance of our method in terms of global interpretability?
The RQ3 involves three objectives: (1) evaluating the global interpretability of our approach, (2) assessing its computational efficiency, and (3) demonstrating its reliability. For the first objective, we aim to demonstrate that our approach provides decision tree-like interpretability. The decision tree is a typical global interpretation method that represents the overall decision logic through hierarchical “IF-Then” rules [45]. Furthermore, we adopt the same quantitative method as [46] to evaluate global interpretability by measuring the contributions of metric subsets, identified by the interpreter, to the model’s predictive performance. Specifically, we use Joint_SDT and SP-LIME to generate six metric subsets for each dataset and use them to retrain Joint_CNN independently. The predictive performance of Joint_CNN with different metric subsets is measured by AUC. Higher AUC values indicate that the interpreter provides better global interpretability. For the second objective, we assess the computational efficiency of our method and SP-Lime by comparing the training times. Since excessive training time is a key limitation of existing global interpretation methods [47]. For the third objective, we use three indicators, CD, FI and AI to quantify the reliability of our approach.
RQ4: Do the metrics selected by Joint_SDT have a significant impact on the performance of Joint_CNN?
This study evaluates the consistency between the metrics selected by Joint_SDT and those impact the performance of Joint_CNN most. To address this, we conduct ablation experiments in which each metric identified by Join_SDT is individually removed, and the resulting degradation in Joint_CNN’s performance (e.g., declines in Matthews Correlation Coefficient (MCC), F-measure, and AUC) is measured. MCC is commonly used to assess the quality of imbalanced binary classification, where the value ranges from -1 to 1. -1 indicates complete disagreement between predictions and ground truth, 1 represents perfect agreement. A strong correlation between Joint_SDT’s prioritized metrics and the magnitude of performance deterioration would further confirm that our joint-learning framework make positive contributions to both predictor and interpreter.
$MCC=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (18)
(a) The first interpretation
(b) The second interpretation
(c) The third interpretation
(d) The fourth interpretation
Figure 6: Multiple interpretations of LIME for a certain instance in the PC2 dataset.
IV-A Dataset
Twenty datasets used in the experiment are collected from the AEEEM 111 https://zenodo.org/record/3362613#.YmuGoNpByUk, NASA  [^2] http://promise.site.uottawa.ca/SERepository/datasets-page.html and PROMISE  [^3] http://promise.site.uottawa.ca/SERepository repositories. Numerous studies have demonstrated the efficacy of these datasets [48, 49, 50, 51]. Further details about each dataset are provided in Table II. We observe that these datasets exhibit significant class imbalance, with the highest imbalance rate reaching 0.41% from the PC2 dataset.
TABLE II: Details of the datasets used in the experiment
Group Dataset Instances Defects %Defects
AEEEM JDT 997 206 20.66%
LC 691 64 9.26%
ML 1862 245 13.16%
PDE 1492 209 14.01%
NASA kc1 2109 326 15.46%
mc1 9466 68 0.72%
jm1 7782 1672 21.49%
PC1 549 77 14.03%
PC2 5589 23 0.41%
PC3 1563 160 10.24%
PC4 1458 178 12.21%
PROMISE ant-1.7 745 166 22.28%
camel-1.6 965 188 19.48%
ivy-1.2 352 40 11.36%
jedit-4.1 312 79 25.32%
log4j-1.0 135 34 25.19%
lucene-2.4 340 203 59.71%
poi-3.0 442 281 63.57%
synapse1.2 256 86 33.59%
xerces-1.3 453 69 15.23%
IV-B Setting
We config our model as follows: the learning rate is initialized as 1e-06 and the batch size is set to 16. SGD optimizer is employed with a momentum of 0.9. The depth of SDT is set to 4. The penalty strength of SDT is set to 1e+1 and the penalty decay rate is set to 0.25. The hyperparameters $\alpha$ and $\beta$ in Eq. (2) are set to 1.4 and 0.6, and $\gamma$, $\lambda$ in Eq. (2) are set to 0.8 and 0.5. The temperature hyperparameter T in Eq. (6) is set to 100. The window size for the Exponential Moving Average (EMA) is set to 1000. Training stops when the model reaches the maximum number of iterations (40) or when early stopping criterias are met (no validation loss improvement for 10 consecutive epochs). Based on preliminary experimental results, the maximum number of iterations is set to 40, beyond which model performance saturates with minimal further improvement. The dataset was split into 7:2:1 for training, testing, and validation. The experiments are implemented by TensorFlow 2.3.0 with the NVIDIA GeForce RTX 3090 GPU. To ensure fairness, we applied this configuration to the experiments on all datasets.
IV-B1 Baseline
For RQ1, we utilize LIME [12] and BreakDown [30] as baselines for the following reasons: (1) These methods are extensively used in the SDP research; (2) LIME, BreakDown, and our approach belong to feature-oriented interpretation techniques that aim to quantify the correlations between input features and predictive outputs; (3) The theoretical foundations of LIME and BreakDown are different, thereby enriching the diversity of comparisons. For RQ2, we use SVM, Random Forest (RF), Deep Belief Network (DBN), and AutoSpearman [52] as baselines. These models are selected based on the following reasons: (1) SVM and RF are widely adopted non-explainable SDP models, while DBN represents DL-based approach in SDP. (3) AutoSpearman, in contrast, is a well-recognized explainable SDP model. Including it as a baseline ensures a fair comparison. For RQ3, we use SP-LIME [45] as the baseline for two key reasons: (1) SP-LIME and our method share the same technical foundation. Both of them are perturbation-based global interpretable methods. Moreover, SP-LIME is the representative method in this category [45]. (2) SP-LIME is the global version of Lime (one of the baselines of RQ1). Selecting it as the baseline is helpful to maintain the consistency across the experiments.
IV-B2 Data preprocessing
It involves normalization and data sampling in this section. Specifically, normalization is defined as follows:
$\widetilde{x}=\frac{x-\min{\left(D\right)}}{\max{\left(D\right)}-\min{\left(D\right)}}$ (19)
where $x$ is the sample of a dataset, and ${\rm min}(\cdot),{\rm max}(\cdot)$ are the minimum and maximum operations applied column-wise across the dataset. Normalization ensures that all features are on the same scale, and it helps the algorithm converge faster and more smoothly. To address the class imbalance issue, we adopt two sampling techniques: SMOTE (Synthetic Minority Oversampling Technique) [3] and RUS (Random Undersampling) [53].
IV-B3 Statistical significance test
To verify the significance of the improvements, we introduce $Cohen^{\prime}s\ d$ in this section, which takes any value between 0 and infinity. When $Cohen^{\prime}s\ d$ takes values in the ranges $0,0.2),[0.2,0.8),and\ [0.8,+\infty)$, it indicates that the improvement is small, medium, and large 444 https://en.wikipedia.org/wiki/Effect_size. The definition of $Cohen^{\prime}s\ d$ is as follows:
$Cohen^{\prime}s\ d=\frac{M_{1}-M_{2}}{\sqrt{(SD_{1}^{2}+SD_{2}^{2})/2}}$ (20)
where $M_{1}$ and $M_{2}$ represent the means of the two group data, and $SD_{1}$ and $SD_{2}$ represent their standard deviations.
TABLE III: The CD-10%s of LIME, BreakDown and proposed method
LIME BreakDown Joint_SDT
Average CD-10% 65% 74.5% 100%
Improvement 53.85%↑ 34.23%↑ -
TABLE IV: The AI and FI of Joint_SDT and LIME
AIAI FIFI
LIME Joint_SDT LIME Joint_SDT
JDT 0.8200 0.8200 0.8100 0.9300
LC 0.8143 0.8429 0.8000 0.8714
ML 0.6684 0.9037 0.6791 0.9786
PDE 0.7200 0.7533 0.6533 0.7867
jm1 0.6239 0.8010 0.8793 0.9910
kc1 0.6540 0.8531 0.5355 0.9431
mc1 0.8754 0.9916 0.9155 0.9979
PC1 0.8727 0.8727 0.7091 0.9273
PC2 0.9821 0.9982 0.9714 0.9803
PC3 0.7834 0.9172 0.8471 0.9682
PC4 0.8151 0.9178 0.9041 0.9932
ant-1.7 0.6667 0.7467 0.6667 0.9200
camel1.6 0.7010 0.5361 0.7010 0.8144
ivy-1.2 0.8056 0.8333 0.8056 0.9167
jedit-4.1 0.5312 0.7500 0.5312 0.5938
log4j-1.0 0.7857 0.6429 0.7857 0.5714
lucene-2.4 0.5000 0.7059 0.5000 0.7059
poi-3.0 0.4222 0.7778 0.4222 0.8222
synapse-1.2 0.5769 0.6538 0.5769 0.5769
xerces-1.3 0.6522 0.7174 0.6522 0.8478
Average 0.7135 0.8018 0.7173 0.8568
Improvement 12.38%↑ - 19.45%↑ -
$Cohen^{\prime}s\ d$ 0.6792 - 0.9451 -
TABLE V: Prediction Performance of Joint_CNN and Base_CNN in Terms of F-measure, AUC, PPC, PNPC and FOR
F-measure AUC PPC PNPC FOR
Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN
JDT 0.6802 0.8544 0.7297 0.8360 0.6000 0.6900 0.4000 0.3100 0.2683 0.1585
LC 0.5686 0.8285 0.6923 0.6862 0.7714 0.7857 0.2286 0.2143 0.1692 0.1538
ML 0.5686 0.7123 0.6897 0.6418 0.6203 0.6952 0.3797 0.3048 0.3216 0.2398
PDE 0.7475 0.7111 0.7039 0.7032 0.7000 0.7867 0.3000 0.2133 0.2105 0.1128
jm1 0.5755 0.8621 0.6240 0.6280 0.5417 0.8036 0.4583 0.1964 0.1335 0.0012
kc1 0.6443 0.7303 0.7545 0.8087 0.5024 0.5308 0.4976 0.4692 0.0783 0.0345
mc1 0.7352 0.8621 0.8082 0.6280 0.8786 0.7818 0.1214 0.2182 0.0012 0.0652
PC1 0.7475 0.8621 0.7039 0.6280 0.6909 0.8182 0.3091 0.1818 0.0500 0.0816
PC2 0.4996 0.8621 0.5000 0.6280 0.9911 0.9875 0.0089 0.0125 0.0018 0.0018
PC3 0.5588 0.7957 0.7634 0.8324 0.7707 0.6752 0.2293 0.3248 0.0397 0.0093
PC4 0.7021 0.7472 0.8169 0.7843 0.8082 0.6644 0.1918 0.3356 0.0167 0.0300
ant-1.7 0.7991 0.7907 0.8276 0.8022 0.7733 0.5733 0.2267 0.4267 0.2267 0.0444
camel-1.6 0.6932 0.7337 0.5992 0.6687 0.5567 0.5876 0.4433 0.4124 0.1692 0.1739
ivy-1.2 0.8273 0.8347 0.6016 0.8594 0.7778 0.7500 0.2222 0.2500 0.0667 0.1000
jedit-4.1 0.5862 0.6946 0.5429 0.8125 0.5312 0.5938 0.4688 0.4062 0.0556 0.0500
log4j-1.0 0.7755 0.6520 0.8500 0.8500 0.5714 0.3571 0.4286 0.6429 0.2727 0.0235
lucene-2.4 0.5623 0.6947 0.5482 0.7643 0.2941 0.3529 0.7059 0.6471 0.4737 0.3684
poi-3.0 0.7156 0.6951 0.7263 0.7414 0.4000 0.5333 0.6000 0.4667 0.3793 0.1724
synapse1.2 0.6616 0.6486 0.8235 0.7516 0.3077 0.6154 0.6923 0.3846 0.2000 0.1111
xerces-1.3 0.7921 0.8407 0.8498 0.7875 0.6304 0.6957 0.3696 0.3043 0.1212 0.1111
Average 0.6720 0.7706 0.7078 0.7421 0.6359 0.6639 0.3641 0.3361 0.1628 0.1022
Improvement 14.67%↑ - 4.85%↑ - 4.4%↑ - 8.33% ↑ - 59.30%↑ -
$Cohen^{\prime}s\ d$ 1.1321 - 0.3553 - 0.0763 - 0.1665 - 0.5338 -
TABLE VI: The F-measure and AUC performance of different models on all datasets
F-measure AUC
Dataset SVM RF DBN AutoSpearman Joint_CNN SVM RF DBN AutoSpearman Joint_CNN
JDT 0.6239 0.6989 0.7162 0.7824 0.8544 0.7554 0.7419 0.7852 0.8598 0.8360
LC 0.3467 0.7177 0.3467 0.7871 0.8285 0.5846 0.8462 0.5846 0.5046 0.6862
ML 0.4919 0.5958 0.3219 0.8054 0.7123 0.6732 0.7522 0.4596 0.7339 0.6418
PDE 0.5565 0.6209 0.4874 0.7837 0.7111 0.6875 0.6601 0.6130 0.6957 0.7032
jm1 0.4476 0.5849 0.4476 0.7580 0.8621 0.5030 0.5762 0.5030 0.6269 0.6280
kc1 0.5695 0.5371 0.6918 0.5814 0.7303 0.6992 0.6141 0.7724 0.8111 0.8087
mc1 0.4754 0.5513 0.4742 0.5814 0.8621 0.8914 0.5577 0.8326 0.9373 0.6280
PC1 0.3870 0.7556 0.3939 0.8276 0.8621 0.6562 0.8557 0.6057 0.6815 0.6280
PC2 0.4736 0.4996 0.4213 0.9819 0.8621 0.4507 0.5000 0.3647 0.4194 0.6280
PC3 0.5580 0.5745 0.5224 0.8345 0.7957 0.7914 0.5772 0.7356 0.8336 0.8324
PC4 0.5700 0.6694 0.3535 0.8617 0.7472 0.7380 0.6154 0.5651 0.7814 0.7843
ant-1.7 0.7900 0.7121 0.5446 0.7760 0.7907 0.8409 0.7477 0.6591 0.8266 0.8022
camel-1.6 0.4920 0.5367 0.6314 0.6710 0.7337 0.6199 0.5621 0.6923 0.6356 0.6687
ivy-1.2 0.6244 0.7317 0.5000 0.8721 0.8347 0.6719 0.8125 0.7500 0.6250 0.8594
jedit-4.1 0.3914 0.3273 0.3431 0.7585 0.6946 0.4657 0.3743 0.4771 0.8698 0.8125
log4j-1.0 0.4286 0.3538 0.4286 0.6533 0.6520 0.6364 0.5909 0.6364 0.7750 0.8500
lucene-2.4 0.7153 0.3585 0.5983 0.7059 0.6947 0.7140 0.5000 0.6333 0.8250 0.7643
poi-3.0 0.6584 0.6786 0.5139 0.7819 0.6951 0.6463 0.6624 0.5392 0.8082 0.7414
synapse-1.2 0.5439 0.6067 0.3210 0.5247 0.6486 0.5437 0.6062 0.4125 0.6732 0.7516
xerces-1.3 0.6430 0.6865 0.4697 0.7611 0.8407 0.7747 0.6758 0.6465 0.7766 0.7875
Average 0.5394 0.5899 0.4764 0.7545 0.7706 0.6690 0.6453 0.6146 0.7350 0.7421
Improvement 42.86%↑ 30.63%↑ 61.75%↑ 2.13%↑ - 10.93%↑ 15.00%↑ 20.75%↑ 0.97%↑ -
$Cohen^{\prime}s\ d$ 2.3697 1.7196 2.9703 0.1712 - 0.7337 0.9543 1.1906 0.0657 -
TABLE VII: The CD, FI and AI values of Joint_SDT and SP-Lime
Method CD FI AI
Joint_SDT 1.00 0.8714 0.8429
SP-Lime 0.81 0.7933 0.8012
IV-C Experimental results
The experimental results are presented as follows.
IV-C1 The results of RQ1
To evaluate the consistency of Joint_SDT, LIME and BreakDown, we applied each of them to interpret the same instance ten times. As shown in Table III, the average CD-10% scores of LIME, BreakDown and Joint_SDT are 65%, 74.5%, and 100%. This result suggests that both LIME and BreakDown exhibit poor consistency. To further investigate the root cause of such poor consistency, we visualized the interpretations of LIME in Fig. 6. We can see that LIME is unable to assign consistent weights to the same metric across multiple explanations, which we believe it is the root of poor consistency. Moreover, we conducted a comparison between the interpretations generated by LIME and Joint_SDT in terms of AI and FI. Since BreakDown does not provide an explicit surrogate function like LIME [12], we are unable to calculate its AI and FI. As shown in Table IV, Joint_SDT achieves an average improvement of 12.38% in AI and 19.45% in FI compared to LIME. Furthermore, the $Cohen^{\prime}s\ d$ results indicate that the significance of improvement in AI is medium, while in FI, it is large. This results further confirm that Joint_SDT outperforms LIME in the consistency of interpretation.
Although Joint_SDT generally shows improvements in AI and FI, its performance is worse than LIME on certain datasets, such as Log4j. This is largely due to the limited amount of data, where sparse distributions enable LIME to maintain consistent interpretability by reducing sampling randomness. However, the neural network-like structure of Joint_SDT requires a larger volume of data for stable and reliable feature importance estimation. As a result, Joint_SDT struggles to achieve desirable performance on small datasets.
IV-C2 The results of RQ2
According to Table V, we observe that Joint_CNN outperforms Base_CNN in terms of F-measure, AUC, PPC, PNPC, and FOR on most datasets, and the average improvements are 14.67%, 4.85%, 4.4%, 8.33% and 59.30% respectively. The corresponding $Cohen^{\prime}s\ d$ test results indicate that the effect sizes of the improvements in the five indicators are large, medium, small, small, and medium. From these results, we can draw the following conclusions: (1) The improvement of Joint_CNN over Base_CNN across all evaluation metrics indicates that the joint-learning framework makes a positive contribution in improving the performance of SDP models. (2) The improvements in PNPC and FOR suggest that Joint_CNN is more reliable in distinguishing defective and non-defective instances, which is crucial for real-world SDP applications. (3) The constant improvements across all datasets suggests that the joint-learning framework generalize well and is not limited to specific dataset.
Furthermore, according to Table VI, joint_CNN outperforms SVM, RF, DBN and AutoSpearman in both F-measure and AUC. Specifically, Joint_CNN achieves average improvements of 42.86%, 30.63%, 61.75% and 2.13% in F-measure and 10.93%, 15%, 20.75% and 0.97% in AUC. The $Cohen^{\prime}s\ d$ test results indicate that the effect sizes of of improvements in F-measure are large, large, large, and small, while those in AUC are medium, large, large, and small. According to these results, we can confirm that Joint_CNN achieves better accuracy and demonstrates strong robustness across all dataset.
We believe there are two reasons related to the improvements: (1) We explicitly integrate the prediction error $L_{pred}(f,D)$ into the loss function of the joint-learning framework, guiding the predictor to capture discriminative information contained in the data. (2) The revised distillation loss $L_{int}$ can be regarded as a penalty term of $L$, which effectively encourages the framework to learn more generalized features of the data, improving the model’s robustness.
IV-C3 The results of RQ3
To demonstrate the global interpretability of our approach, we use the PC1 dataset as an example. As presented in Fig. 5, the decision logic can be abstracted as the path from the “DECISION_DENSITY” metric (denoted as the root) to the leaf node (denoted as 1 or 0). For example, one feasible decision path can be presented as DECISION_DENSITY -> MAINTENANCE_SEVERITY -> PERCENT_COMMENTS -> PERCENT_COMMENTS -> 0, which is similar to the interpretation format of a decision tree. Furthermore, following the framework of [46], we visualized the AUCs of Joint_CNNs using different metric subsets identified by Joint_SDT and SP-LIME in Fig. 7. The results demonstrate that our method achieves higher AUC values. This confirms that the metric subsets selected by Joint_SDT are more effective for explaining the model’s global decision-making logic compared to those selected by SP-LIME.
We measure the computational cost by calculating the average training time. Specifically, we employ the Python $\mathrm{time}$ module to record the start and end times of the training process for both our approach and SP-Lime. To ensure statistical reliability, this procedure is repeated ten times for each method. The results show that the average computational costs for our approach and SP-Lime are 744.7 seconds and 2301.4 seconds, respectively. These findings indicate that our method achieves a significant reduction in computational cost compared to SP-Lime. The substantial difference can be attributed to the optimized design of our method, which abstracts the decision logic as the path of SDT from the root node to the leaf node. In contrast, SP-Lime requires extensive perturbations and evaluations across the feature space, leading to higher computational requirements.
Furthermore, we calculate the CD, FI and AI values for LC dataset to evaluate the reliability of our approach and SP-Lime. As presented in table VII, our method achieves a better performance in all three indicators. The reason for this result can be partially attributed to the random sampling involved in perturbation processes of SP-Lime [45].
Figure 7: Comparison of AUC values of Joint_CNNs using metric subsets selected by Joint_SDT and SP-LIME on LC, PC1, and Lucene datasets.
(a) F-measure
(b) AUC
(c) MCC
Figure 8: Ablation study results across multiple indicators. The gray dashed line indicates the performance of DP-CNN before removing metrics.
TABLE VIII: Rankings of Feature Importance
Dataset The Most Important Feature Identified by Joint_SDT F-measure Rank AUC Rank MCC Rank
JDT WCHU_numberOfMethodsInherited 4 3 5
LC WCHU_cbo 2 1 2
ML ck_oo_numberOfPublicMethods 3 2 5
PDE WCHU_fanOut 3 3 5
jm1 HALSTEAD_DIFFICULTY 3 2 1
kc1 v 1 2 2
mc1 LOC_TOTAL 1 1 1
PC1 branchCount IV 2 1 1
PC2 DECISION_DENSITY 1 4 1
PC3 DESIGN_DENSITY 4 3 3
PC4 LOC_TOTAL 2 2 1
ant-1.7 mfa 1 5 1
camel-1.6 lcom 4 3 4
ivy-1.2 cam 2 5 4
jedit-4.1 lcom3 4 5 4
log4j-1.0 max_cc 4 4 3
lucene-2.4 dam 1 1 1
poi-3.0 wmc 2 3 2
synapse-1.2 cbo 1 1 4
xerces-1.3 max_cc 2 5 2
IV-C4 The results of RQ4
According to the structure of Joint_SDT, the most important metric is positioned at the root of the SDT. For instance, in the PC2 dataset, ”DECISION_DENSITY” is identified as the most important metric. To validate this finding, we conduct an ablation experiment that removes each metric selected by Joint_SDT individually and the corresponding changes are observed in the performance of Joint_CNN. The ranking of performance degradation reflects the importance of each metric to the model. The ablation experiment results for PC2 are visualized in Fig.  8. We can see that removing ”DECISION_DENSITY” results in the most significant performance decline across the three indicators: MCC, F-measure, and AUC. This confirms that ”DECISION_DENSITY” is the most important metric for PC2 dataset. This aligns with the interpretation of our method.
To enhance the comprehensiveness of the analysis, we extended the ablation experiment to all datasets. Table VIII presents the most important metrics identified by Joint_SDT across the twenty dataset. The results show that the root metrics identified by Joint_SDT typically correspond to the most significant performance degradations in the ablation experiments.
V Threats to validity
While the proposed model demonstrates improvements in both prediction accuracy and interpretability, certain limitations may affect its generalizability. The details are as follows.
  • Choice of dataset: The experiments were conducted on the publicly available datasets. These datasets may not fully capture the diversity of software systems in real-world scenarios, potentially limiting the generalizability of the model.
  • Hyperparameter setting: The model’s performance is related with the setting of hyperparameters for both the predictor and interpreter. Suboptimal setting may lead to less accurate predictions or interpretations.
  • Application scenarios: In this paper, we focus on the within-project SDP. However, the performance of our model in the context of the cross-project setting has not been evaluated.
VI Conclusion and future works
In this paper, we explore the possibility of designing a defect predictor and its corresponding interpreter collaboratively. Unlike most existing approaches, which treat defect prediction and interpretation as separate tasks, we regard them as strongly correlated. We introduce a framework aiming at improving the reliability of interpretation and enhancing predictive accuracy simultaneously. Based on the extensive empirical evaluations, we obtained the following key findings:
  • Compared to existing interpretation methods such as LIME and BreakDown, the proposed framework significantly improves the reliability of interpretations through the incorporation of the KD principle.
  • The joint-learning framework demonstrates superior accuracy on widely used datasets by explicitly incorporating interpretation results into the loss function.
Future research will explore the following directions:
  • Using diverse real-world datasets to evaluate the robustness and generalization of our method.
  • Integrating automatic hyperparameter optimization to reduce manual tuning efforts.
  • Assess the model’s performance in different defect prediction scenarios, such as cross-project defect prediction.